Home‎ > ‎Reports‎ > ‎Crew100B‎ > ‎

day10


Crew 100B Log Book for March 7, 2011

Photos of the day





Commander's Check-In Report

Balwant Rai Reporting

Physical Status: Excellent

Experiment Status: All the medical, astrobiological, rover and food experiments began today with unique results.

Time Departed/Returned from EVA: three EVAs performed

The meals today were very delicious since it was a cooking day.

EVA’s: We completed three successful EVA’s.

Number Of reports: 3 science reports, 1 engineering report, 1 journalist report, 3 EVA reports and 1 chef report.

Plan for tomorrow: will initiate the medical experiment again on all crew members. After that, we will do simulated microgravity experiments. Interpretation of the Sun by means of the radio telescope will also commence. The study of the rover’s maneuverability as well as the potential use of the drone, weather pending, will take place. Soil samples will yet again be collected during two EVA’s tomorrow. Tomorrow will be a non-cooking day.


Commander's Report

Balwant Rai Reporting

Today we started the day continuing medical experiments and our habitability experiment.

The crew’s medical health will continue to be monitored by our health and safety officer for the remainder of our stay here on Mars. Breakfast was satisfying. Our crew had their saliva samples taken (via Versi Sal 1 from Oasis Diagnostics), vital parameters monitored (via the Zephyr Bioharness), and their heart rate variability measured (via NERV Express 4.2 software). I also performed the cognitive study (via CogState Research software) on all crew members, as well. We performed an emergency rescue experiment for all crew members to be well aware of their tasks and react in well time to avoid confusions and time waste.

Today all crew members had to go through their dental check up done by our health and safety officer. Our crew scientist continued to work through problems with starting observations of the sun via the radio telescope. Our rover engineer completed her third experiment. Three EVA’s were performed to sites of interest. Today is a cooking day, so everyone enjoyed meals served at lunch and hope the same for dinner.Our astronomer continued her experiments well.

Human machine, simulated microgravity experiments, Medical, habitability and astrobiological experiments will be taken into account, also, the rover experiment and the radio telescope experiments will be performed.


Engineering Report

Matthieu Ansart Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): between 0.75 and 0.5

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.35

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 2 containers full

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage)


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 10

· Trailer Potable Water Tank Level (inches from bottom): Empty

· Hab Potable Water Tank Level (inches from bottom): 9

· Potable Water Meter Reading:

Notes/Comments : No leak


GreenHab:

· Condition of plants in Tank 1: Green water

· Condition of plants in Tank 2: Green plants and snails

· Condition of plants in Tank 3: Green plants

· Greenhab notes/ comments: Tank 2 full


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: No

· Fuel Consumed:Replaced

· Tire Status: Good

· Comments:


ATV 2 (Viking I):

· Used: Yes

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Good

· Comments:


ATV 3 (Viking II):

· Used: Yes

· Oil Checked: No

· Fuel Consumed:Replaced

· Tire Status: Good

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 66

· Actual temperature upstairs (Farenheit): 73

· Thermostat setting downstairs (Farenheit): 55

· Actual temperature downstairs (Farenheit): 61


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering

1. Notes/Comments:


Science Report (Medical Rescue)

Balwant Rai and Jasdeep Kaur Reporting

During long duration space flights such as Mars missions, astronauts are prone to many physiological changes such as loss of bone mass, muscle strength, and cardiovascular fitness. Maintenance of these factors and putting then under control is important requisite; so, we planned the EVA for medical rescue, a crew member with muscular injury of leg during driving ATV was rescued with balm massage and hot bandage.

Lessons Learned: Every ATV should have first aid kit to manage the minor injury .


Science Report (Human Factors)

Balwant Rai Reporting

Leadership issues with multicultural crews on MDRS : Lessons learned from MDRS Crew 100 B ILEWG Euro MoonMars crew

In isolated and confined environments such as in MDRS, Utah ( MDRS Crew 100B ILEWG Euro MoonMars crew) , two important leadership roles have been identified: the task oriented role and the supportive role. On the space missions, the commander should be familiar with both of these aspects of leadership and to explain the same to crew members.In this study, I found for crew-members that the supportive role of the commander and solving problem while discussing after lunch and dinner could make a huge difference in social behavior of each member towards each other.To solve the existing problems collectively in free time leads to crew cohesion and motivations. For mission control personnel on the ground, both the task and supportive roles of their leader is very important to build crew member as a team to have additional output and good results.


Science Report (Oral Hygiene)

Jasdeep Kaur Reporting

Time: 0930-dental examination was done for all crew members

Crew Members: Science Team: 6 crew members Balwant Rai, Rachel Dompnier , Quentin Bourges,Crystal Latham,Matthieu Ansart and Jasdeep Kaur

Site Location: Hab

Objective: To figure out dental findings and other oral findings after one week of stay in Hab. PREVENTION IS BETTER THAN CURE

Observation: All crew members were observed for their dental findings and oral findings. All teeth were examined for caries, periodontal infections, gums, saliva, tongue, mucosa, TMJ problems. Mouth opening was noted for all members once again.

Results: As with respect to the principles to be followed on Mars, crew members cannot use non-biodegradable products, so toothpastes, floss, mouthwashes cannot be used and crew members have to follow the rules to brush the teeth with baking soda. It was observed that all crew members developed little plaque on teeth due to poor oral hygiene maintenance. As the crew members are not allowed to use toothpastes and mouthwashes, so warm saline measures were suggested for all members. All crew members had little low mouth opening as compared to previous dental check up, due to little environment change that could have affected the stress markers.

Brushing Technique: Warm saline rinses were explained to all crew members for proper maintenance of oral health.


EVA 29 Report

Quentin Bourges Reporting

EVA#29 09:30 am to 10:30 pm

Crew members: Rachel Dompnier, Matthieu Ansart, Quentin Bourges

Location: in the vicinity of the Hab

Transit mode: Foot

Objective: Cooperation/Coordination with Rover and the UAV.

Results: This morning’s experiment was a failure as we had problem with Wi-Fi connection. We couldn’t get good response from Rover’s commands.

Lesson learned: None, as the mission was continually interrupted by Wi-Fi connection problems.


EVA 30 Report

Crystal Latham Reporting

EVA #30: 14:22 - 15:39

Crew members: Matthieu Ansart, Quentin Bourgès, Crystal Latham

Location: To the North about 2.4 km away from the HAB

Transit mode: ATV/foot

Objective: To take a panoramic picture of a specific location, as well as a 360° high-frame rate series of pictures of a specific area.

Results: We were able to successfully take these photos.

Lessons learned: When searching for a specific location with the GPS you should follow in the direction opposite which the GPS arrow is pointing, to find said location. Wearing any type of glasses is highly suggested when driving ATVs in cold/ muddy conditions.


EVA 31 Report

Rachel Dompnier Reporting

EVA #31, 4:45 to 5:10 pm

Crew Members: Jasdeep Kaur, Balwant Rai, Rachel Dompnier

Site Location: N38.42887 W110.79982

Transit Mode: ATV’s

Objectives: To practice the emergency procedure for someone who was injured on ATV

Results: We provided first medical care and carefully came back to the Hab.

Lesson learned: As per science report


Chef’s Report

Jasdeep Kaur Reporting

1a) Today was a cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

No

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Tofu with spicy sweet-corns and peas, chicken cream soup, spinach bread and veggie-fruity salad

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

No

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

No

5) Lessons learned (if any)

No

6) Comments/questions for Kim and Jean, the food study investigators

Nothing


Journalist Report

Jasdeep Kaur Reporting

It is crew 100B’s tenth full day on Mars. Today all crew members had to go through their dental check-up once again and they were advised to use warm saline rinses. Our medical tests are carried out the same way everyday.

It started raining in the morning, but it did not last long and crew could again plan to go out for ATVs. A few crew members participated in the first EVA of the day and rest of them participated the next one after that. Our crew astronomer has been trying to get the radio telescope operational, with no avail so far. It was really nice weather after rain stopped and EVAs were really mood altering.

A unique EVA for rescue was done on ATVs. One of crew member pretended to hit a lump of mud and was diverted away to hit another lump(enacted). The crew members got muscles stretched and could not raise the leg any longer. So, first aid measures were given and crew member rescued. Today was a cooking day, so everybody enjoyed meals served at lunch and hope the same for dinner. Tonight, we plan to discuss with all crew members about their likes and dislikes at MDRS. Also, we will have a short discussion of our output for this mission.